OpenStudy (viixxen):

What are the coordinates of the vertex of the graph of the function y = -3x^2 - 12x + 3? a.) (-2, 29) b.) (2, -15) c.) (2, -9) d.) (-2, 15) I think it might be B, but I'm not sure.

OpenStudy (anonymous):

did you draw the graph

OpenStudy (btaylor):

the x-value of the graph of a parabola, form ax^2+bx+c is -b/2a. If you then plug that x-value in, you get the y-value of the vertex.

OpenStudy (btaylor):

here, -b/2a=-(-12)/2(-3) --> 12/-6 = -2 y evaluated at x=-2 gives: y=-3(4)-12(-2)+3=-12+24+3=15 So, your answer is (-2, 15). It is D; you were close, though.

OpenStudy (anonymous):

I think its -2,15 because if im not wrong it has to be -2

OpenStudy (smored):

change it to vertex form. \[y= -3x ^{2}-12x+3\] \[y=-3(x ^{2}+4x-1)\] \[y=-3[(x ^{2}+4x+4)-1-4]\] \[y=-3[(x+2) ^{2}-5]\] \[y=-3(x+2) ^{2}+15\] so vertex is (-2,15)

OpenStudy (smored):

btw, \[y=a(x-h) ^{2}+k\] where vertex is (h,k)

OpenStudy (anonymous):

I don't think I learned the vertex form !!! I hate teachers who don't know how to teach !!!

OpenStudy (smored):

oh if you did not, you will later. you can use \[-\frac{ b }{ 2a }\] to find the x coordinate of the vertex, the find the y coordinate by inserting the answer for x.

OpenStudy (smored):

that was what I learned before the vertex form in algebra.

OpenStudy (btaylor):

so exactly what I said...?

OpenStudy (smored):

yep

OpenStudy (smored):

you should use the Equation form, it is easier to read.

OpenStudy (btaylor):

I know. I don't feel like taking the time sometimes.